# references.bib

### v2

@@ Abstract
Multi-scale systems often exhibit stochastic and deterministic dynamics. Capturing these aspects in a compartmental model is challenging. Notably, low occupancy compartments exhibit stochastic dynamics and high occupancy compartments exhibit deterministic dynamics. Failing to account for stochasticity in small populations can produce ‘atto-foxes’, e.g. in the Lotka-Volterra ordinary differential equation (ODE) model. This limitation becomes problematic when studying extinction of species or the clearance of infection, but it can be resolved by using discrete stochastic models e.g. continuous time Markov chains (CTMCs). Unfortunately, simulating CTMCs is infeasible for most realistic populations.
We develop a novel mathematical framework, to couple continuous ODEs and discrete CTMCs: ‘Jump-Switch-Flow’ (JSF). In this framework populations can reach extinct states (“absorbing states”), thereby resolving atto-fox-type problems. JSF has the desired behaviours of exact CTMC simulation, but is substantially faster than existing alternatives.
JSF’s utility for simulation-based inference, particularly multi-scale problems, is demonstrated by several case-studies. In a simulation study, we demonstrate how JSF can enable a more nuanced analysis of the efficacy of public health interventions. We also carry out a novel analysis of longitudinal within-host data from SARS-CoV-2 infections to quantify the timing of viral clearance. JSF offers a novel approach to compartmental model development and simulation.
@@
Dynamical systems represent a powerful method for describing the world and have successfully been applied in many fields. However, modelling populations which change in size across multiple scales remains a challenge. Population sizes in biological processes often change over orders of magnitude, e.g. the spread of infectious disease (anderson1991infectious), the boom-and-bust of insect populations (ludwig1978qualitative), and the immune response to infections including HIV (perelson2002modelling), and influenza virus (baccam2006kinetics).
@@
The dynamics of small populations can be heavily influenced by stochastic effects, while for larger populations these fluctuations often average out, justifying the use of continuum models. When the population under consideration does not change in size over orders of magnitude, there is usually a natural (and obvious) choice between a stochastic or deterministic model. However, for multi-scale models, this remains a challenge. Developing modelling methods that can bridge these scales has been a long-standing goal of applied mathematics (cotter2016error; flegg2014analysis; isaacson2013convergent).
@@
Compartmental models describe how quantities (e.g. the number of cells, people, or molecules) change in a dynamical system. Two popular ways to represent compartmental models are ordinary differential equations (ODEs) and continuous time Markov chains (CTMCs). Many ODEs one encounters are actually “ensemble averages” of CTMCs, although the mathematical justification of this is not always straightforward (Kurtz_1970; Kurtz_1971; kurtz1972relationship).
@@
In ODEs, the state of the system changes continuously. This can lead to “atto-fox problems” where populations shrink to infeasibly small sizes where a real population would have gone extinct (fowler2021atto; mollison1991dependence; lobry2015migrations). In CTMCs, the discrete state changes stochastically and there may be absorbing states e.g. extinction of a population. CTMCs offer a more natural description of small populations sizes, however applying them to large populations can be challenging. Moreover, a range of powerful techniques can be brought to bear on ODEs, enabling more thorough analysis. This raises a natural question about what to do when compartment occupancy changes across orders of magnitude.
@@
There are exact and approximate algorithms to simulate CTMCs simoni2019stochastic. Exact methods (e.g. Doob-Gillespie, first/next reaction, and rejection based methods) are computationally expensive when state transitions occur at a high rate (sanft2015constant). Approximate methods (e.g. Tau-leaping (gillespie2001approximate) and chemical Langevin equations (rao2003stochastic; cao2006efficient; gillespie2000chemical; gibson2000efficient)) scale to high transition rates but can have unacceptable approximation error. To overcome the limitations of classical approximate methods, over the past two decades, there has been a concerted effort to develop hybrid stochastic-deterministic approaches (simoni2019stochastic).
@@
Some of the approaches to hybrid modelling involve partitioning the transitions in CTMC into fast and slow transitions which can be sampled individually and subsequently synchronised is efficient but complex to implement simoni2019stochastic. Jump-diffusion differential equations partition the model compartments into fluid and discrete to construct hybrid switching jump diffusion processes buckwar2011rungekutta; angius2015approximate. Recently, kynaston2023regime considered extended systems that represent each compartment with both a continuous and a discrete version and allow the for conversion between them.
@@
Essential for the real-world utility of any mathematical model of a dynamical system is its ability to be calibrated to data. For deterministic compartment models, standard parameter inference methods such as maximum likelihood and posterior sampling via Markov chain Monte Carlo (MCMC) can readily be applied (ionides2006inference). For stochastic compartment models these inference problems are usually much harder. Simulation-based inference has emerged as a way to handle these problems, e.g. particle filter (arulampalam2002tutorial; kitagawa1996monte) and approximate Bayesian computation  (Alahmadi2020ABC).
@@
In this paper, we present a simple and efficient hybrid simulation method: Jump-Switch-Flow (JSF). This method enables adaptive transitions between stochastic and deterministic regimes across compartments, ensuring conservation principles are maintained. In this approach, compartments can individually switch regimes, allowing for the compartments to be split into stochastic and deterministic subsets, while still being coupled. We demonstrate the utility and properties of the JSF method through simulation studies and a case study in which we analyse existing longitudinal SARS-CoV-2 viral load dataset.
@@
Two simulation studies are presented to explore the properties of JSF and demonstrate its use in an inference setting. By comparing simulations from an epidemic model, we contrast the computational efficiency and accuracy of JSF with the exact Doob-Gillespie method and Tau-leaping. A simulation study demonstrates the types of insight available when using an approach supporting absorbing (extinction) states.
@@
The significance of being able to do inference with models with absorbing states, i.e. models in which one of the compartments can go extinct, is demonstrated by a case study in which we analyse longitudinal SARS-CoV-2 viral load data ke2022daily using a TEIRV (Target cells – Eclipsed cell – Infectious cells – Refractory cell – Virions) model to infer the state of host and virus across the infection. Understanding how a viral infection is cleared has important ramifications for treatment and prevention. For example, the initial exposure may fail to initiate a systematic infection, understanding the conditions under which this happens is important for infection prevention pearson2011stochastic; the infection may reach low levels, potentially escaping detection, but the virus may rebound and cause disease, understanding when the virus has been cleared is important for understanding how long treatment must be maintained. Inferring viral clearance can be computationally challenging yan2016extinction, this case study demonstrates how JSF enables the tractable estimation of viral clearance.
@@ Methods
Consider a compartmental model with $n$ compartments $\vec{V}=\left\{V_{i}\right\}_{i=1}^{n}$ where $V_{i}(t)$ represents the value of the $i$th compartment at time $t$. For example, $V_{i}$ could be the number of people infected with a pathogen, or the copy number of a molecule in a cell. The state variables $V_{i}$ may take values from different domains depending upon the resolution needed for the model. For example, in an ODE, $\vec{V}$ will have real values and in a CTMC $\vec{V}$ might have integer values.
@@
Typically, discrete values are used to represent small populations, while larger populations will be represented with a continuum. To accommodate both scales, we model the domain of $V_{i}$ as $\mathcal{V}_{\Omega_{i}}=\{0,1,\ldots,\Omega_{i}\}\cup(\Omega_{i},\infty)$. The switching threshold parameter, $\Omega_{i}\in\mathbb{Z}_{\geq 0}$, is where the $i$th compartment transitions from discrete to continuous dynamics. If a compartment $V_{i}$ has a value in $\{0,1,\ldots,\Omega_{i}\}$, we call it discrete (or jumping), and if it has a value in $(\Omega_{i},\infty)$, we call it continuous (or flowing). While the switching threshold can be compartment specific, for ease of exposition, we will only consider a single threshold shared between all compartments $\Omega=\Omega_{i}$. At any moment in time let us assume $q$ of the $n$ compartments are flowing. We use the notation $\vec{V}_{F}=\left\{V_{i}:V_{i}>\Omega\right\}\in(\Omega,\infty)^{q}$ and $\vec{V}_{J}=\left\{V_{i}:V_{i}\leq\Omega\right\}\in\left\{0,1,\ldots,\Omega\right\}^{(n-q)}$ to represent the compartments in each of the flowing and jumping states, respectively.
@@
The dynamics of each compartment $V_{i}$ are described by a set of $m$ reactions =$\mathcal{R}=\left\{\mathcal{R}_{k}\right\}_{k=1}^{m}$. Each reaction $\mathcal{R}_{k}$ is defined by two properties: the rate (per unit time) at which it occurs, $\lambda_{k}$, which may be (and usually is) a function of the state $\vec{V}$; and the effect on the state, i.e. the change $\eta_{ik}$ to the size of compartment $V_{i}$ when reaction $\mathcal{R}_{k}$ occurs. As a matrix, $\eta\in\mathbb{Z}^{n,m}$ is referred to as the stoichiometric matrix. For ODE models, these reactions occur continuously and are written in the form
$\frac{\mathrm{d}\vec{V}}{\mathrm{d}t}=\eta\vec{\lambda}(\vec{V}),$ (1)
@@
while for CTMC models, reactions in the system $\mathcal{R}$ occur as discrete events. In the later case, each reaction $\mathcal{R}_{k}$ has a separate propensity described by $\lambda_{k}(\vec{V})$, this propensity remains constant between events but when an event $\mathcal{R}_{k}$ occurs, there is a change in $\vec{V}$ (as specified by the elements of $\eta_{\cdot k}$) and therefore in $\vec{\lambda}(\vec{V})$.
@@
Consider for example the SIR model. The state of this model is $\vec{V}=(S,I,R)^{\intercal}$, and there are two “reactions”: infections ($\mathcal{R}_{1}$) and recoveries ($\mathcal{R}_{2}$). For infections, the rate of reaction is $\lambda_{1}=\beta V_{1}V_{2}/(V_{1}+V_{2}+V_{3})$, and the entries of the stoichiometric matrix are $\eta_{1,1}=-1$, $\eta_{2,1}=1$ and $\eta_{3,1}=0$. For recoveries, the rate of reaction is $\lambda_{2}=\gamma V_{2}$, and the entries of the stoichiometric matrix are, $\eta_{1,2}=0$, $\eta_{2,2}=-1$, and $\eta_{3,2}=1$.
@@
We define the subset of reactions $\mathcal{S}\subseteq\mathcal{R}$ to contain those treated as stochastic events. The set $\mathcal{S}$ can be defined in one of two ways. We only use the second in this manuscript, but describe both to assist the presentation. The first way to define $\mathcal{S}$ is as follows; $\mathcal{S}=\left\{\mathcal{R}_{k}:\exists i\text{ s.t. }V_{i}\in\vec{V}_{J}\text{ and }\eta_{ik}\neq 0\right\}$. In this definition, a reaction $\mathcal{R}_{k}$ is included in $\mathcal{S}$ if the reaction has an effect on a jumping (discrete) population $V_{i}\in\vec{V}_{J}$. This is a minimum requirement; a reaction should not be permitted to evoke a continuous change in a discrete population. However, we do want to allow reactions to make discrete changes to flowing (continuous) populations. The second way to define $\mathcal{S}$, which we use throughout this manuscript, captures a larger set of reactions; $\mathcal{S}=\left\{\mathcal{R}_{k}:\exists i\text{ s.t. }V_{i}\in\vec{V}_{J}\text{ and }\left(\eta_{ik}\neq 0\text{ or }\partial_{V_{i}}\lambda_{k}\neq 0\right)\right\}$. In this definition, a reaction is included in $\mathcal{S}$ if either (1) it causes a change in jumping (discrete) populations or (2) it is influenced by a discrete population (perhaps as reactants for example).
@@
Reactions in $\mathcal{S}$ are simulated using stochastically sampled times similar to CTMC models (SI Section 1A). It is important to note that, unlike time homogeneous CTMC models, the propensities are not constant because the state $\vec{V}$ (and therefore $\vec{\lambda}$) are continuously varying. When any reaction $\mathcal{R}_{k}\in\mathcal{S}$ occurs, we say the system has jumped and an instantaneous change of $\eta_{ik}$ for each compartment $V_{i}$ occurs (irrespective of whether $V_{i}\in\vec{V}_{J}$ or $V_{i}\in\vec{V}_{F}$ to ensure mass conservation is observed). We will refer therefore to reactions in $\mathcal{S}$ as jumps. The reactions in $\mathcal{S}^{\prime}=\mathcal{R}\setminus\mathcal{S}$ are not stochastic, we call these flows because they represent the continual change of value of the relevant compartments, all of which are continuous by definition of $\mathcal{S}^{\prime}$ (SI Section 1B). At any moment in time, we denote $|\mathcal{S}^{\prime}|=p=m-|\mathcal{S}|$ to be the number of reactions which are flowing.
@@
Finally, the hybrid model that we propose is capable of switching. Switch events are defined as a compartment between $\vec{V}_{F}$ and $\vec{V}_{J}$. These events occur when a compartment’s value crosses the switching threshold $\Omega$ (SI Sections 1C and 1D). Importantly, switch events can change $\mathcal{S}$ and are paradigm defining events which should occur infrequently compared to jumps (frequent) and flows (continuous).
@@
Due to the way that $\mathcal{R}$ is partitioned, it is possible to order the rows and columns of $\eta$ at any moment into the upper-triangular block form
$\eta=\left(\begin{array}[]{c|c}\eta_{\mathcal{S}^{\prime}}&\bar{\eta}_{\mathcal{S}}\\ \hline\cr 0&\eta_{\mathcal{S}}\end{array}\right),$
@@
where $\eta_{\mathcal{S}^{\prime}}\in\mathbb{Z}^{q\times p}$, $\eta_{\mathcal{S}}\in\mathbb{Z}^{(n-q)\times(m-p)}$ and $\bar{\eta}_{\mathcal{S}}\in\mathbb{Z}^{q\times(m-p)}$ refer to stoichiometric coefficients for changes in flowing compartments under flows, jumping compartments under jumps, and flowing compartments under jumps, respectively. Written as a system of equations analogous to (1), the hybrid JSF model we propose formally takes the following form. For any time interval $t_{0}<t<t_{1}$ between switching events,
$\frac{\mathrm{d}\vec{V}_{F}}{\mathrm{d}t}$ $=\eta_{\mathcal{S}^{\prime}}\vec{\lambda}_{\mathcal{S}^{\prime}}(\vec{V})+\bar{\eta}_{\mathcal{S}}\vec{\Lambda}_{\mathcal{S}}(\vec{V}),$
$\vec{V}_{J}(t)$ $=\vec{V}_{J}(t_{0})+\eta_{\mathcal{S}}\int_{t_{0}}^{t}\vec{\Lambda}_{\mathcal{S}}(\vec{V})\ \mathrm{d}s,$
@@
where $\vec{\lambda}_{\mathcal{S}^{\prime}}\in\mathbb{R}^{p}$ are the reaction rates of flows and $\vec{\Lambda}_{\mathcal{S}}$ is a stochastic vector of $m-p$ delta-function spike trains that are derived from the realisations of $m-p$ different jumps sampled at rates which are dependent on the dynamic changes in the propensities $\vec{\lambda}_{\mathcal{S}}\in\mathbb{R}^{m-p}$ for these jumps (SI Section 1A).
@@
The Lotka-Volterra model describes the populations of two species: prey, $V_{1}$, and predators, $V_{2}$. The prey reproduce at a rate $\alpha$, the predators die at a rate $\gamma$ and new predators are produced through predation at rate $\beta$. When modelled with ODEs, this gives us the following system:
@@
Fig. 1A shows a compartmental diagram of the Lotka-Volterra model with the reaction: $\mathcal{R}_{1}$, birth of $V_{1}$; $\mathcal{R}_{2}$, death of $V_{2}$; and $\mathcal{R}_{3}$, conversion of $V_{1}$ into $V_{2}$. Fig. 1B shows a representation of the reactions and their representation with a stoichiometric matrix.
Fig. 1C shows a cartoon trajectory of this process when represented with JSF. There are three switching events (at $t_{1}$, $t_{2}$ and $t_{3}$), and each configuration of discrete and continuously changing state features:

### v3

# A hybrid framework for compartmental models enabling simulation-based inference

@@ Abstract
Multi-scale systems often exhibit a combination of stochastic and deterministic dynamics. In compartmental models, low occupancy compartments tend to exhibit stochastic dynamics while high occupancy compartments tend to follow deterministic dynamics. Representing both dynamics with existing methods is challenging. Failing to account for stochasticity in small populations can produce “atto-foxes”, for example in the Lotka-Volterra ordinary differential equation (ODE) model. This limitation becomes problematic when studying the extinction of species or the clearance of infection, but it can be overcome by using discrete stochastic models, such as continuous time Markov chains (CTMCs). Unfortunately, simulating CTMCs is impractical for many realistic models, where discrete events have very high frequencies.
In this work, we develop a novel mathematical framework to couple continuous ODEs and discrete CTMCs: “Jump-Switch-Flow” (JSF). In this framework, compartments can reach extinct states (“absorbing states”), thereby resolving atto-fox-type problems. JSF has the desired behaviours of exact CTMC simulation, but is substantially computationally faster than existing alternatives, by at least one order of magnitude, and can even obtain constant scaling, irrespective of compartment occupancy.
We demonstrate JSF’s utility for simulation-based inference, particularly multi-scale problems, with several case-studies. In a simulation study, we demonstrate how JSF can enable a more nuanced analysis of the efficacy of public health interventions. We also carry out a novel analysis of longitudinal within-host data from SARS-CoV-2 infections to quantify the timing of viral clearance. In this work, we show how JSF offers a novel approach to compartmental model simulation.
@@ Introduction
Dynamical systems are a powerful method for describing the world, and have successfully been applied in many fields. However, modelling populations which change in size across multiple scales remains a challenge [14]. Population sizes in biological processes often change over orders of magnitude, e.g. the spread of infectious disease [2]; the boom-and-bust of insect populations [32]; and the immune response to infections, such as HIV [38], and influenza virus [5].
@@
The dynamics of small populations can be heavily influenced by stochastic effects, while for larger populations these fluctuations often average out, justifying the use of continuum models. When the population under consideration does not change in size over orders of magnitude, there is usually a natural (and obvious) choice between a stochastic or deterministic model. However, for multi-scale models, this decision remains a challenge, and developing modelling methods that can bridge these scales has been a long-standing goal of applied mathematics [10, 13, 22].
@@
Compartmental models describe how quantities (e.g. the number of molecule, cells, or people) change in a dynamical system. Two popular ways to represent compartmental models are ordinary differential equations (ODEs) and continuous time Markov chains (CTMCs). Many ODEs one encounters are actually ensemble averages of CTMCs, although the mathematical justification of this is not always straightforward [27, 28, 29].
@@
In ODEs, the state of the system changes continuously, which can lead to “atto-fox problems” where populations shrink to infeasibly small sizes, i.e. where a real population would likely have gone extinct [14, 34, 31]. In CTMCs, the discrete state changes stochastically and there may be absorbing states e.g. extinction of a population. As a result, CTMCs offer a more natural description of small populations sizes, however applying them to large populations can be computationally challenging. Moreover, a range of powerful techniques can be brought to bear on ODEs, enabling more thorough analysis. This raises a natural question about what to do when compartment occupancy changes across orders of magnitude.
@@
Exact and approximate algorithms to simulate CTMCs exist [16, 42]. Exact methods (e.g. Doob-Gillespie, first/next reaction, and rejection based methods) are computationally expensive when state transitions occur at a high rate [41]. Approximate methods (e.g. Tau-leaping [18, 9] and the use of chemical Langevin equations [39, 17, 15]) scale to high transition rates but can have unacceptable approximation error. To overcome the limitations of classical approximate methods, over the past two decades, there has been a concerted effort to develop hybrid stochastic-deterministic approaches [42, 26, 7, 40].
@@
Some hybrid modelling approaches involve partitioning the transitions in a CTMC into fast and slow transitions. Despite having significant implementation bookkeeping and overheads, these approaches can be sampled individually and subsequently synchronised in an efficient manner [42]. Jump-diffusion differential equations partition the model compartments into fluid and discrete compartments, to construct hybrid switching jump diffusion processes [8, 3]. Other hybrid simulation techniques have been developed in the context of within-host viral infection [26]. However, this implementation is not generalised, but tailored to a single system, and does not allow switching between stochastic and deterministic regimes. In the context of determining the hybridisation between stochastic and deterministic couplings, path-integrals have been utilised, [7], which permit analysis of hybrid processes. Recently, [30] considered extended systems that represent each compartment with both a continuous and a discrete version and allow the for conversion between them [30].
@@
Essential for the real-world utility of many mathematical models is their ability to be calibrated to data. For deterministic compartment models, standard parameter inference methods, such as maximum likelihood and posterior sampling via Markov chain Monte Carlo (MCMC), can readily be applied [21]. For stochastic models these inference problems are usually much harder. Simulation-based inference has emerged as a way to handle these problems, e.g. the particle filter [4, 24] and approximate Bayesian computation [1].
Calibrated multi-scale simulations of ODE (compartment) models is an important milestone towards accurate predictive models that address stochastic behaviour. In a recent paper, Trindade and Zygalakis develop a hybrid scheme for chemical kinetics over different scales with the goal of demonstrating the utility for parameter estimation  [43]. In this paper, a CTMC is used to model reaction events on small populations (below a threshold $I_{1}>0$) and for better efficiency Tau-leaping is used for reactions that only impact large populations (above a threshold $I_{2}>I_{1}$). Between these two thresholds, a “blending function” is used to create a linear transition between an accurate CTMC and efficient Tau-leaping treatment of the reactions. The hybrid adoption of CTMC and Tau-leaping regimes has the property of maintaining appropriate levels of noise at large populations but struggles as populations grow much larger where a Chemical Langevin or deterministic ODE description may be more computationally appropriate.
@@
In this paper, we present a simple and efficient hybrid simulation method which we call Jump-Switch-Flow (JSF). This method enables adaptive transitions between stochastic and deterministic regimes across compartments, while ensuring conservation principles are maintained. In this approach, compartments can individually switch regimes, allowing for the compartments to be split into stochastic and deterministic subsets, while remaining coupled. This regime switching presents unique challenges in matching at interfaces: ensuring that the deterministic regime represents the expected state under transition between regimes and that mass is conserved. Conservation is not always essential when it pertains to small population sizes, however in many cases sustained small populations have a significant impact on larger populations (e.g. enzymes in chemical processes and viruses within hosts.) We demonstrate the utility and properties of the JSF method through simulation studies, and a case study in which we reanalyse existing longitudinal SARS-CoV-2 viral load datasets.
@@
Two simulation studies are presented to explore the properties of JSF and demonstrate its use in an inference setting. In the first simulation study, we compare the computational efficiency and accuracy of JSF with the exact Doob-Gillespie method and the Tau-leaping method. In the second simulation study, we demonstrate the types of insight available when using an approach that supports absorbing (extinction) states.
@@
In the case study, we demonstrate The significance of being able to perform inference with models with absorbing states (i.e. models in which one of the compartments can go extinct.) In the case study we reanalyse longitudinal SARS-CoV-2 viral load data [23] using JSF to simulate a TEIRV (Target cells – Eclipsed cell – Infectious cells – Refractory cell – Virions) model and infer the state of host and virus across the infection. Understanding how a viral infection is cleared has important ramifications for both treatment and prevention. For example, the initial exposure may fail to initiate a systematic infection, and understanding the conditions under which this happens is important for infection prevention [37]. Moreover, the infection may reach low levels, potentially escaping detection, but the virus may rebound and cause disease. Understanding when the virus has been cleared is important for optimising treatment duration. Inferring viral clearance can be computationally challenging [44], and this case study demonstrates how JSF enables the tractable estimation of viral clearance.
@@ Methods
Consider a compartmental model with $n$ compartments, $\vec{V}=\left(v_{1},...,v_{n}\right)$, where the state variable, $v_{i}:=v_{i}(t)$, represents the value of the $i$th compartment at time $t$. For example, $v_{i}$ could be the number of people infected with a pathogen, or the copy number of a molecule in a cell. The state variables $v_{i}$ may take values from different domains, depending upon the resolution needed, for example in an ODE, $v_{i}$ will have real values and in a CTMC $v_{i}$ might have integer values.
@@
When the variables represent quantities, typically, discrete values are used to represent small populations, while larger populations are represented with a continuum. To accommodate both scales, let the domain of $v_{i}$ be the set $\{0,1,\ldots,\Omega_{i}\}\cup(\Omega_{i},\infty)$. The switching threshold parameter, $\Omega_{i}\in\mathbb{Z}_{\geq 0}$, is where the $i$th compartment transitions between discrete and continuous dynamics. If a compartment $v_{i}\in\{0,1,\ldots,\Omega_{i}\}$, we call it discrete (or jumping), and if $v_{i}\in(\Omega_{i},\infty)$, we call it continuous (or flowing). While the switching threshold can be compartment specific, for ease of exposition, we will only consider a single threshold shared between all compartments, i.e. $\Omega_{i}=\Omega$.
At each time $t$, the components of $\vec{V}$ can be partitioned into $\vec{V}_{F}$ and $\vec{V}_{J}$, where $\vec{V}_{F}$ contains the flowing variables $v_{i}>\Omega$, and $\vec{V}_{J}$ contains the jumping variables $v_{i}\leq\Omega$. Therefore, at any moment in time $\dim{(\vec{V}_{F})}$ of the $n$ compartments are flowing, and $\dim{(\vec{V}_{J})}$ of the $n$ compartments are jumping, with $\dim{(\vec{V}_{F})}+\dim{(\vec{V}_{J})}=n$.
@@
The dynamics of each compartment $v_{i}$ are described by a set of $m$ reactions =$\mathcal{R}=\left\{{r}_{k}\right\}_{k=1}^{m}$. Each reaction ${r}_{k}$ is defined by two properties: the rate (per unit time) at which it occurs, $\lambda_{k}$, which (usually) is a function of the state $\vec{V}$; and the reaction’s effect on the state, i.e. the change $\eta_{ik}$ to the size of compartment $v_{i}$ when reaction ${r}_{k}$ occurs. As a vector, $\vec{\lambda}\in\mathbb{R}^{m}$ is referred to as the propensity vector. As a matrix, $\eta\in\mathbb{Z}^{n,m}$ is referred to as the stoichiometric matrix. For ODE models that only contain flowing variables, these reactions occur continuously and are written in the form:
$\frac{\mathrm{d}\vec{V}}{\mathrm{d}t}=\eta\vec{\lambda}(t,\vec{V}).$ (1)
@@
For CTMC models, reactions in the system $\mathcal{R}$ occur as discrete events. In the latter case, each reaction ${r}_{k}$ has a separate propensity described by $\lambda_{k}(t,\vec{V})$. This propensity remains constant between reactions, but when a reaction ${r}_{k}$ occurs, there is a change in $\vec{V}$ (as specified by the elements of $\eta_{\cdot k}$), and therefore in $\vec{\lambda}(t,\vec{V})$.
@@
As an example, consider the SIR model of epidemics, which, in ordinary differential equations, has the form $\frac{dS}{dt}=-\frac{\beta SI}{S+I+R}$, $\frac{dI}{dt}=\frac{\beta SI}{S+I+R}-\gamma I$, and $\frac{dR}{dt}=\gamma I$. The state of this model is $\vec{V}=(S,I,R)^{\intercal}$, and there are two “reactions”: infections (${r}_{1}$) and recoveries (${r}_{2}$). For infections, the rate of reaction may be modelled by $\lambda_{1}=\beta SI/(S+I+R)$, and the entries of the associated column of the stoichiometric matrix are $\eta_{1,1}=-1$, $\eta_{2,1}=1$ and $\eta_{3,1}=0$. For recoveries, the rate of reaction may be modelled by $\lambda_{2}=\gamma I$, and the entries of the associated column of the stoichiometric matrix are, $\eta_{1,2}=0$, $\eta_{2,2}=-1$, and $\eta_{3,2}=1$.
@@
We define $\mathcal{R}_{J}\subseteq\mathcal{R}$ as the subset of reactions treated as stochastic jump events. The set $\mathcal{R}_{J}$ can be defined more precisely in two ways. While we use only the second in this manuscript, both are described for clarity. The first way to define $\mathcal{R}_{J}$ is as: $\mathcal{R}_{J}=\left\{{r}_{k}:\exists i\text{ s.t. }v_{i}\in\vec{V}_{J}\text{ and }\eta_{ik}\neq 0\right\}$. In this definition, a reaction ${r}_{k}$ is included in $\mathcal{R}_{J}$ if and only if the reaction has some material effect on a jumping (discrete) compartment $v_{i}\in\vec{V}_{J}$. This is a minimum requirement; a reaction should not be permitted to evoke a continuous change in a discrete compartment. However, where it makes sense to do so, it is possible to allow reactions to make discrete changes to flowing (continuous) compartments. The second way to define $\mathcal{R}_{J}$, which we use throughout this manuscript, captures a larger set of reactions: $\mathcal{R}_{J}=\left\{{r}_{k}:\exists i\text{ s.t. }v_{i}\in\vec{V}_{J}\text{ and }\left(\eta_{ik}\neq 0\text{ or }\partial_{v_{i}}\lambda_{k}\neq 0\right)\right\}$. In this definition, a reaction is included in $\mathcal{R}_{J}$ if either (1) it causes a material change in jumping (discrete) compartments or (2) it is influenced/caused by a discrete compartment (e.g. as catalyst of some reaction.)
@@
Reactions in $\mathcal{R}_{J}$ are simulated using stochastically sampled times, similar to CTMC models (and described below in the section on ‘Jump events’). It is important to note that unlike time homogeneous CTMC models, the propensities are not constant because the state $\vec{V}$ (and therefore $\vec{\lambda}$) are continuously varying. When any reaction ${r}_{k}\in\mathcal{R}_{J}$ occurs, we say the system has jumped and an instantaneous change of $\eta_{ik}$ for each compartment $v_{i}$ occurs (irrespective of whether $v_{i}\in\vec{V}_{J}$ or $v_{i}\in\vec{V}_{F}$, to ensure the conservation of mass). We therefore refer to reactions in $\mathcal{R}_{J}$ as jumps. The reactions in $\mathcal{R}_{F}=\mathcal{R}\setminus\mathcal{R}_{J}$ represent the continuous change of value of the relevant compartments, all of which are deterministic (see section on ‘Flow events’ above.) We therefore refer to reactions in $\mathcal{R}_{F}$ as flows. At any moment in time $|\mathcal{R}_{F}|$ is the number of reactions which are flowing, and $|\mathcal{R}_{J}|$ is the number of reactions which are jumping, such that $|\mathcal{R}_{F}|+|\mathcal{R}_{J}|=m$.
@@
Finally, the hybrid model that we propose is capable of switching. Switching events occur when the size of a compartment $v_{i}$ crosses the threshold $\Omega$ and therefore changes membership between the flowing $\vec{V}_{F}$ and jumping $\vec{V}_{J}$ sets (see sections on ‘Jump clock updates’ and ‘Switching events’). Importantly, switching events may also impact which reactions are jumps, $\mathcal{R}_{J}$. As such, switches are paradigm-defining events which ought to occur infrequently in comparison to jumps (which may occur frequently) and flows (which occur continuously).
@@
Due to the way that $\mathcal{R}$ is partitioned, it is possible to order the rows and columns of $\eta$ into the upper-triangular block form:
$\eta=\left(\begin{array}[]{c|c}\eta_{FF}&\eta_{FJ}\\ \hline\cr 0&\eta_{JJ}\end{array}\right),$ (4)
@@
where $\eta_{FF}\in\mathbb{Z}^{\dim{(\vec{V}_{F})},|\mathcal{R}_{F}|}$, $\eta_{JJ}\in\mathbb{Z}^{\dim{(\vec{V}_{J})},|\mathcal{R}_{J}|}$ and ${\eta}_{FJ}\in\mathbb{Z}^{\dim{(\vec{V}_{F})},|\mathcal{R}_{J}|}$ refer to stoichiometric coefficients for changes in flowing compartments under flows, jumping compartments under jumps, and flowing compartments under jumps, respectively. Note that, the lower left corner is zeros because, by definition, flowing reactions cannot influence the state of jumping variables. Written as a system of equations analogous to (1), the hybrid JSF framework we propose formally takes the following form. For any time interval $t_{i}<t<t_{i+1}$ between switching events $i$ and $i+1$:
$\frac{\mathrm{d}\vec{V}_{F}}{\mathrm{d}t}$ $=\eta_{FF}\vec{\lambda}_{F}(t,\vec{V})+{\eta}_{FJ}\vec{\Lambda}_{J}(t,\vec{V}),$ (5)
$\vec{V}_{J}(t)$ $=\vec{V}_{J}(t_{i})+\eta_{JJ}\int_{t_{i}}^{t}\vec{\Lambda}_{J}(s,\vec{V})\ \mathrm{d}s,$ (6)
@@
where $\vec{\lambda}_{F}\in\mathbb{R}^{|\mathcal{R}_{F}|}$ are the reaction rates of flows and $\vec{\Lambda}_{J}$ is a stochastic vector of $|\mathcal{R}_{J}|$ delta-function spike trains that are derived from the realisations of $|\mathcal{R}_{J}|$ different jumps sampled at rates which are dependent on the dynamic changes in the propensities $\vec{\lambda}_{J}\in\mathbb{R}^{|\mathcal{R}_{J}|}$ for these jumps (see ‘Jump events’).
Figure 1: JSF provides a way to capture both the continuous deterministic dynamics of large populations and the discrete stochastic dynamics of small populations. A A compartmental model representation of the Lotka-Volterra system (7): prey ($v_{1}$, depicted as red rabbits which, in reaction $r_{1}$, reproduce at rate $\alpha$), predators ($v_{2}$, depicted as purple foxes which, in reaction $r_{2}$, die at rate $\gamma$), and their interaction (predation of rabbits but foxes, in reaction $r_{3}$, at a rate proportional to $\beta$). B The compartmental model can be formalised as a reaction network via a stoichiometry matrix $\eta$ which details each species role in the reactions. The reactants consumed $\eta^{-}$ and the products produced $\eta^{+}$ can be written explicitly with $\eta=\eta^{+}-\eta^{-}$. C A hypothetical JSF trajectory demonstrating how the representation captures the continuous variation of some compartment-reaction pairs (flow) and the discrete stochastic changes (jumps) in others. Transitioning (switching) between a continuous and discrete state occurs at a threshold $\Omega$ which is indicated with a horizontal dashed line. D An example trajectory from the Lotka-Volterra model exhibiting the characteristic cyclical behaviour with additional stochasticity influencing the dynamics at low population sizes. E An example trajectory showing the possibility for the predator species to go extinct and the subsequent growth of the prey species.
@@
The Lotka-Volterra model describes the populations of two species: prey, $v_{1}$, and predators, $v_{2}$. The prey reproduce at rate $\alpha$, the predators die at rate $\gamma$ and new predators are produced through predation at rate $\beta$. When modelled with ODEs, this gives us the following system:
@@
Figure 1A shows a compartmental diagram of the Lotka-Volterra model with the reaction: ${r}_{1}$, birth of $v_{1}$; ${r}_{2}$, death of $v_{2}$; and ${r}_{3}$, conversion of $v_{1}$ into $v_{2}$. Figure 1B shows a representation of the reactions and their corresponding stoichiometric matrix.
Figure 1C depicts a hypothetical trajectory of the Lotka-Volterra model when represented as a JSF process. There are three switching events (at $t_{1}$, $t_{2}$ and $t_{3}$), and each configuration of discrete (jumping) and continuous (flowing) states are featured: